OpenStudy (anonymous):

Steve has these scores on the exams for this semester: 89, 92, and 87. What score will he have to get on the fourth exam to have a final average of at least 90%? 1)98% or better 2)96% or better 3)92% or better 4)90% or better

OpenStudy (anonymous):

3

OpenStudy (anonymous):

89 + 92 + 87 + x = 90(4)

OpenStudy (anonymous):

im confused! which one works?

OpenStudy (anonymous):

89+92+87+x=360 268+x=360 x=360-268 x=92

OpenStudy (anonymous):

gotcha! thankyou!